### v4

# A Text-guided Protein Design Framework

Abstract
Current AI-assisted protein design mainly utilizes protein sequential and structural information. Meanwhile, there exists tremendous knowledge curated by humans in the text format describing proteins’ high-level functionalities. Yet, whether the incorporation of such text data can help protein design tasks has not been explored. To bridge this gap, we propose ProteinDT, a multi-modal framework that leverages textual descriptions for protein design. ProteinDT consists of three subsequent steps: ProteinCLAP which aligns the representation of two modalities, a facilitator that generates the protein representation from the text modality, and a decoder that creates the protein sequences from the representation. To train ProteinDT, we construct a large dataset, SwissProtCLAP, with 441K text and protein pairs. We quantitatively verify the effectiveness of ProteinDT on three challenging tasks: (1) over 90% accuracy for text-guided protein generation; (2) best hit ratio on 12 zero-shot text-guided protein editing tasks; (3) superior performance on four out of six protein property prediction benchmarks.
Machine learning (ML) has recently shown profound potential for protein discovery. These ML tools have been quickly adapted as auxiliary and accelerating roles in scientific pipelines, including but not limited to protein engineering [1], structure prediction [2], structure reconstruction [3], and inverse folding [4].
To this end, most existing such methods merely model the intrinsic properties of protein sequences and structures [5, 6, 7, 8, 9, 10]. Motivated by recent breakthroughs of foundation models [11, 12, 13, 14], approaches in the computational chemistry domain that utilize textual descriptions of drugs in addition to their intrinsic chemical and structural information have proved to be effective in small-molecule drug discovery [15, 16, 17, 18, 19]. This naturally raises a question for the protein domain: how can such a multi-modal mechanism benefit protein engineering and generation? To answer this question, we initiate a novel paradigm that leverages textual descriptions for protein design.
Specifically, we aim to exploit the following two modalities of proteins: the protein sequence and the textual description. The protein sequence is the sequence of 20 types of amino acids (a.k.a. residues) that influence how the protein will fold and what functions it can perform; while the second modality is the textual description recorded in the open data sources, e.g., UniProt [20]. Such annotation data can contain abundant knowledge of proteins [21], such as biological processes they participate in, molecular functions they execute, and cellular components they localize to. These two modalities respectively focus on expressing the internal biochemical composition and the high-level knowledge summarized by domain experts. Thus, it is worth exploring the incorporation of both to realize more challenging and promising protein design tasks, such as zero-shot generalization.
Our Approach. To attain the aforementioned goal, we propose Protein Design with Text (ProteinDT), a multi-modal framework for protein design. ProteinDT is composed of three key steps, as shown in Figure 1. (1) A Contrastive LAnguage and Protein pretraining (ProteinCLAP) step to align the representations between text sequences and protein sequences. ProteinCLAP adopts the contrastive learning paradigm to align the representation space of the two modalities. This is enabled by a constructed dataset called SwissProtCLAP, with around 441K text-protein pairs extracted from the SwissProt subset of UniProt. (2) A ProteinFacilitator model that produces the protein sequence representation from the text modality. We leverage a Gaussian distribution to estimate this conditional distribution. (3) A decoder model for protein generation. It is a conditional generative model that generates protein sequences conditioned on protein information, i.e., the representation produced by the previous step. Concretely, we present a robust solution to the demanding task by introducing the most powerful generative models for sequence data: one autoregressive (AR) and two diffusion models (ProteinDiff).
We consider three downstream tasks to verify the versatile functionalities of ProteinDT. The first task is text-to-protein generation, where we follow the ProteinDT pipeline to generate protein sequences from the input text prompts, which describe the desired protein attributes. Using AR and ProteinDiff, the optimal retrieval accuracy can reach over 90%. The second task is zero-shot text-guided protein editing, where the inputs are text prompts and protein sequences. We introduce two editing methods: latent interpolation and latent optimization. Latent interpolation conducts an interpolation in the sequence-level representation space, while latent optimization directly optimizes the token-level representation. Both editing methods inject the text modality and the learned representation is used for protein generation accordingly. We observe that ProteinDT can reach the overall best hit ratio on 12 editing tasks, ranging from structure- and stability-aware tasks, to peptide-binding editing tasks. Additional qualitative results offer further compelling evidence supporting the validity of ProteinDT. The last task is protein property prediction, aiming to evaluate the robustness and generalization ability of the learned representation from ProteinCLAP. Compared to six state-of-the-art protein sequence representation methods, ProteinDT can obtain superior performance on four of six benchmark tasks.
Figure 1: Pipeline for ProteinDT pretraining framework (a-c) and downstream tasks (d-f). (a) ProteinCLAP, a contrastive learning paradigm, aligns the representation space of the text and protein sequence modalities. (b) ProteinFacilitator model augments the mapping from text sequence representation to protein sequence representation. (c) A protein sequence decoder, which generates protein sequences conditioned on the representations produced from previous steps. (d) Downstream text-to-protein generation task. (e) Downstream text-guided protein editing task. (f) Downstream protein property prediction task.
Results
In this section, we first provide the preliminaries of ProteinDT and dataset construction. Then we assess three downstream tasks to verify the versatile functionalities of ProteinDT. We focus on exploring two attributes of ProteinDT: (1) ProteinDT can enable zero-shot protein design and editing based on textual descriptions, and (2) ProteinDT can learn a robust protein representation. The specific descriptions for the three downstream tasks, model implementation details, and training hyperparameters can be found in Supplementary B and C.
Dataset Construction and Two Modalities
We construct SwissProtCLAP, a text-protein pair dataset, from UniProt [20]. UniProt is the most comprehensive knowledge base for protein sequences. It also contains extensive annotations including function descriptions, domains, subcellular localization, post-translational modifications, and functionally characterized variants [22]. SwissProt [23], originally created in 1986, is now the subset of UniProt that contains manually reviewed information and expert curation.
Based on SwissProt, we collect and curate a dataset called SwissProtCLAP, with 441,013 text-protein sequence pairs in total. The 441K proteins can be categorized into 327,577 genes and 13,339 organisms. One concrete example can be found in Figure 1, and more details are in Supplementary B.
We also note that the dataset size of SwissProtCLAP is relatively small compared to the peers from other domains (e.g., 400M text-image pairs in the language-vision domain [11]). To alleviate the data insufficiency issue, instead of training our model from scratch, we adopt pretrained representation encoders to train ProteinDT, as elaborated below.
Protein sequence ${\bm{x}}_{\text{p}}$ and representation ${\bm{z}}_{\text{p}}$. Proteins have four levels of structure [24]: primary, secondary, tertiary, and quaternary structure, and the corresponding degree of complexity in the polypeptide chain increases. In this work, we are interested in the protein primary structure ${\bm{x}}_{\text{p}}$, i.e., the amino acid sequence. There are in total 20 different amino acids in proteins, represented with 20 characters, as shown in Figure 1. For the protein sequence-level representation ${\bm{z}}_{\text{p}}=f_{\text{p}}({\bm{x}}_{\text{p}})$, we use BERT as the encoder [25]. It is one of the most advanced Transformer models for modeling sequential data [26]. Additionally, we take the BERT model with ProtBERT pretraining [6]: it was pretrained on a large protein sequence corpus [27, 28], and the pretraining task was to reconstruct the masked tokens.
Text sequence ${\bm{x}}_{\text{t}}$ and representation ${\bm{z}}_{\text{t}}$. The textual descriptions ${\bm{x}}_{\text{t}}$ are formulated in the form of sequences of tokens, and we also adopt the BERT model as the encoder to get the sequence-level representation ${\bm{z}}_{\text{t}}=f_{\text{t}}({\bm{x}}_{\text{t}})$. We further adapt the pretrained SciBERT [29], which was pretrained on the full-text computer science and biomedical papers from Semantic Scholar [30]. Notice that although both modalities use BERT-like models, the detailed architectures and token vocabulary are different. Further, such a disentangled design enables us to adopt the respective pretrained models (e.g., ProtBERT and SciBERT) to mitigate the limited availability of structured text-protein sequence pairs for training.
Pretraining: ProteinDT
The core idea of ProteinDT is the integration of textual descriptions and protein sequences. Such jointly learned knowledge empowers ML models to form informative protein representations with perception for both modalities. As a result, these representations not only contain text-centric domain knowledge about proteins, but also enable the tuning of the language descriptions to accomplish the text-guided protein generation and optimization tasks.
As illustrated in Figure 1, through leveraging the constructed SwissProtCLAP dataset, ProteinDT attains the aforementioned goal with three key steps: a pretraining model, a facilitator, and a decoder. First, the ProteinCLAP model in ProteinDT utilizes contrastive learning to construct representations that integrate knowledge from both text and protein sequences. By doing so, ProteinCLAP not only forms an informative representation of proteins, but also enables the text-protein representation alignment, enabling rich cross-modality tasks like text-protein retrieval. The second step is to train a ProteinFacilitator model, which is an augmented alignment module in addition to ProteinCLAP. To be more concrete, the ProteinFacilitator model maps a piece of text prompt (e.g., “A protein that is stable.") to a corresponding protein representation that captures both the semantic information of the given text and essential protein sequence patterns. Last but not least, ProteinDT subsequently utilizes a conditional generative model to decode the formed protein representation into the protein sequence. We propose one autoregressive model (AR) and two diffusion models (ProteinDiff). More details of the three pretraining steps can be found in the Methods section.
Downstream Task: Text-to-Protein Generation
To illustrate that natural language can guide protein semantics through ProteinDT, we propose the text-to-protein generation task. The pipeline is shown in Figure 1. For the decoder, we consider a Transformer-based AR model and two ProteinDiff models (RNN and Transformer as the transition network, respectively). Notice that we have two options for the condition in decoding. (1) Without ProteinFacilitator: For the input text sequence, we initially obtain the text sequence representation using ProteinCLAP. Subsequently, we directly utilize this representation as a condition to generate the protein sequence using the pretrained encoder. (2) With ProteinFacilitator: Alternatively we encode the input text sequence and employ ProteinFacilitator to derive the protein representation. This modified protein representation is then employed as a condition for generating the protein sequence. An illustration is in Figure 1.
Figure 2: Visualization of text-to-protein generation and text-guided protein editing. (a) Visualization for evaluation of text-to-protein generation. The pretrained ProteinCLAP is used to calculate the similarity between the sampled text and generated protein sequence pairs. (b-c) Two methods for text-guided protein editing: latent interpolation and latent optimization. (d-e) Visualization for evaluation of text-guided protein editing. For four types of editing tasks, different evaluation metrics (marked in red) are applied accordingly.
Table 1: Retrieval accuracy (%) for text-to-protein generation. It is a multiple-choice task ($T$ options), measuring the consistency between the input text prompt and the generated protein sequence.
Galactica ChatGPT ProteinDT (Ours)
w/o ProteinFacilitator w/ ProteinFacilitator
AR AR AR ProteinDiff AR ProteinDiff
Transformer RNN Transformer Transformer RNN Transformer
T = 4 51.5 38.5 49.00 24.00 35.50 97.00 40.50 51.50
T = 10 29.0 23.0 27.00 10.50 17.50 91.00 21.50 25.00
T = 20 19.0 15.5 20.00 5.50 9.50 83.50 15.00 13.50
Experiment setting. We randomly sample 200 text sequences from SwissProtCLAP as the text prompts and then generate a protein sequence for each following the ProteinCLAP-ProteinFacilitator-decoder pipeline and ProteinCLAP-decoder pipeline. For evaluation, we adopt the well-trained ProteinCLAP to obtain the retrieval accuracy: each protein sequence is compared to the input text prompt and $T-1$ randomly chosen texts using ProteinCLAP, and a match is successful if the actual text-sequence pair ranks the highest among all comparisons. The evaluation pipeline is shown in Figure 2, and please check Supplementary D for more details.
Baselines. We consider two large language models as baselines. Galactica (1.3b) [31] is specifically pretrained for scientific tasks. ChatGPT (3.5-Turbo-0301) is one of the most general-purpose foundation models, and existing works have shown its potential in drug editing tasks [19, 32, 33].
Observations. The main results are presented in Table 1, revealing three key observations. Observation 1. Adopting the ProteinFacilitator model is deemed essential. As shown in Table 1, across all ProteinDT decoder models, incorporating ProteinFacilitator yields higher accuracies compared to models without it. Especially for decoders like AR-Transformer, the performance gaps are substantial by up to around 50%. Observation 2. For the three decoder variants of ProteinDT, AR-Transformer can beat the two diffusion models (ProteinDiff). Similar findings are also reported in recent studies on text-to-text generation [34, 35]. We speculate that either the AR model excels in fitting discrete and sequential data models, or we haven’t found the most expressive diffusion model for such a challenging task. These are interesting ML research questions for future exploration. Observation 3. When compared to baseline models, the AR-Transformer with ProteinFacilitator reaches consistently better performance, while the two variants of ProteinDiff with ProteinFacilitator obtain similar performance than Galactica and ChatGPT. It is noteworthy that both baselines are Transformer-based autoregressive models.
Downstream Task: Zero-shot Text-guided Protein Editing
The next task we consider is zero-shot text-guided protein editing. The inputs are an initial protein sequence and a text prompt describing the target property, and the output is the edited protein sequence. The textual modality enables ProteinDT to adapt to flexible target tasks for editing in a zero-shot manner, as long as the text prompt is available. Next, we introduce two text-guided protein editing methods: latent interpolation and latent optimization.
Latent interpolation. Latent interpolation aims at getting the latent codes by directly interpolating two endpoint representations [12, 13, 36], where the latent codes will be used for sampling data points through the decoder. In this setting, we have two representations for the text prompt and input protein as ${\bm{z}}_{\text{t}}$ and ${\bm{z}}_{\text{p}}$, respectively. Then we apply the spherical linear interpolation (SLERP) [13], which yields an intermediate representation $\tilde{\bm{z}}_{\text{c}}=\text{SLERP}({\bm{z}}_{\text{t}},{\bm{z}}_{\text{p}},\theta)$, and the output protein is generated from decoder, as $\tilde{\bm{x}}_{\text{p}}\sim p({\bm{x}}_{\text{p}}|\tilde{\bm{z}}_{\text{c}})$. Here $\theta\in[0,1]$ is a coefficient controlling how close $\tilde{\bm{x}}_{\text{p}}$ is to the text prompt or the input protein, and the larger value of $\theta$ expects $\tilde{\bm{x}}_{\text{p}}$ to be closer to the text prompt.
Latent optimization. We also propose using latent optimization [14, 37, 18] for text-guided protein editing. The high-level idea is to directly optimize a latent code in the representation space and then decode it to get the sampled data. Specifically in ProteinDT, it learns a token-level latent code ${\bm{z}}_{w}$ that is close to both the text and protein sequence-level representation, as ${\bm{z}}_{w}=\arg\min(\lambda\|P({\bm{z}}_{w})-{\bm{z}}_{\text{t}}\|^{2}+(1-\lambda)\|P({\bm{z}}_{w})-{\bm{z}}_{\text{p}}\|^{2})$, where $P$ is the pooling function, and $\lambda$ controls the closeness to the target prompt. Then, the learned ${\bm{x}}_{w}$ are mapped to the protein sequence using a pretrained decoder $g$, i.e., $\tilde{\bm{x}}_{\text{p}}=g({\bm{z}}_{w})$. Notice that $g$ is a token-wise decoder, and can be trained in an auto-encoder manner with the fixed ProteinCLAP protein encoder. More details of the two methods are in Supplementary C.
Editing tasks and evaluation metrics. To fully verify the effectiveness of ProteinDT, we design 12 zero-shot text-guided protein editing tasks. We further categorize them into four settings as detailed below. For editing evaluation, we employ the satisfactory hit ratio (%), which measures the ratio of the output score being higher or lower than the input score, as indicated in the text prompt. The scores are provided by evaluators on each protein sequence, as discussed below. To mitigate bias resulting from different evaluators, we adopt two distinct types of evaluators for each task (except region editing). Certain evaluations require proteins’ 3D structures, for which we employ the ESMFold algorithm [7]. The pipeline of the task design and evaluation are in Figure 2.
(1) Structure editing. Our initial focus involves four editing tasks associated with the secondary structure [19]. The prompts are in the format of “modify the amino acid sequence to have more (less) alpha helices” and “modify the amino acid sequence to have more (less) beta sheets”. We adopt two methods to measure the secondary structures for each amino acid: (1) A protein sequence evaluator for secondary structure prediction (as will be discussed in the property prediction downstream below); (2) Define Secondary Structure of Proteins (DSSP) algorithm for assigning secondary structure to the amino acids [38].
(2) Region editing. The second editing task is region editing. The predicted Local Distance Difference Test (pLDDT) is a per-residue confidence metric used to estimate the reliability of the predicted 3D structure of a protein sequence. We consider proteins from this paper [39], which reveals the relation between ordered/disordered regions and pLDDT. Ordered regions most frequently show higher pLDDT, while the disordered regions are reflected with low pLDDT. The prompts are in the format of “modify the amino acid sequence to have more ordered/disordered regions”. We adopt the pLDDT from ESMFold as the evaluation metric [7].
(3) Stability editing. Then we consider four types of thermodynamic stability editing tasks. Two proteins are considered [40]: villin HP35 and Pin1 WW domain. The prompts are in the format of “modify the amino acid sequence to have higher (lower) stability”. We consider two evaluators: (1) A protein sequence evaluator for stability (as will be discussed in the property prediction downstream below); (2) Rosetta to quantify the energy after structure prediction, in Rosetta Energy Units (REU) [41, 42]. The score function is REF2015 [43].
(4) Peptide binding editing. The last two tasks are peptide binding editing. Peptides are small proteins (usually less than 15 amino acids) with specific functionalities through binding to the target proteins. Here we take 154 pairs of peptide-protein binding pairs from Protein Data Bank (PDB) [44], where the proteins have textual descriptions from UniProt in SwissProtCLAP. Then, the task is to make the input peptide have higher/lower binding affinity with respect to the target protein (textual description). For scoring with the evaluator, we consider two options: (1) A pretrained deep learning model for binding affinity prediction [45]; (2) GNINA to calculate the docking score before and after editing [46].
Baselines. As in the text-to-protein generation tasks, we use Galactica (1.3b) and ChatGPT (3.5-Turbo-0301) for editing baselines. The prompts for ProteinDT and baselines may be modified accordingly. Please refer to Supplementary D for detailed prompts.
Observations. The main results are in Table 2. We can observe that two variants of ProteinDT (latent interpolation and latent optimization) generally perform better on 12 tasks and 21 evaluation methods. However, for the stability editing tasks, latent interpolation methods tend to favor lower stability, while latent optimization can reach a good balance between higher and lower stability tasks. This suggests that latent optimization is, in general, a preferable method for protein editing. Despite being a generic foundation model, ChatGPT is occasionally the second-best performer on some of the stability editing tasks.
We further provide qualitative results in Figure 3. Figure 3 (a) and (d) show the structures of two input protein sequences. Then ProteinDT adopts the prompts with keywords “more $\alpha$-helices” and “more $\beta$-sheets”, the edited protein structures (edited sequences folded with ESMFold) are shown in Figure 3 (b) and (e). Similarly for Figure 3 (c) and (f) when using prompts like “less $\alpha$-helices” and “less $\beta$-sheets”. For peptide editing illustration, we take PDB 3IQI [47], Haemophilus influenzae O-acetylserine sulfhydrylase in complex with the peptide MNENI, as an example. The input peptide is shown in Figure 3 (g), and the output peptides with higher and lower binding affinity are shown in Figure 3 (h, i). We marked the hydrogen bonds in dashed yellow lines, and there are 6, 7, and 2 bonds in Figure 3 (g, h, i), respectively.
Table 2: The satisfactory hit ratio (%) of text-guided protein editing tasks. The best results and the second-best results are marked.
Structure: $\alpha$-helices Structure: $\beta$-sheets Region more less more less ordered disordered BERT DSSP BERT DSSP BERT DSSP BERT DSSP pLDDT pLDDT Galactica 5.46 7.02 13.26 10.53 1.95 2.14 11.50 11.31 8.47 23.73 ChatGPT 5.46 6.82 7.80 6.43 4.68 2.92 5.07 3.51 6.78 62.71 ProteinDT (Ours) Latent Interpolation ProteinDiff-RNN 67.64 75.63 19.49 10.14 27.88 11.31 53.61 69.98 5.08 94.92 ProteinDiff-BERT 38.79 51.27 48.15 34.70 39.38 15.79 42.30 65.50 3.39 94.92 AR-BERT 27.49 45.61 56.92 39.38 9.16 5.26 69.79 73.10 15.25 84.75 Latent Optimization 51.66 54.19 33.92 32.94 46.59 39.57 39.38 46.59 6.78 91.53 Stability: Villin Stability: Pin1 binding editing more stable less stable more stable less stable higher lower BERT Rosetta BERT Rosetta BERT Rosetta BERT Rosetta BindingNet GNINA BindingNet GNINA Galactica 7.81 51.56 73.44 31.25 1.41 4.23 8.45 4.23 26.61 21.77 33.87 25.00 ChatGPT 29.69 57.81 78.12 28.12 28.17 19.72 39.44 71.83 16.94 12.90 26.61 35.48 ProteinDT (Ours) Latent Interpolation ProteinDiff-RNN 17.19 50.00 82.81 54.69 4.23 26.76 95.77 70.42 58.87 29.84 37.90 70.16 ProteinDiff-BERT 9.38 23.44 93.75 68.75 2.82 16.90 97.18 84.51 50.81 29.03 43.55 67.74 AR-BERT 6.25 40.62 93.75 59.38 1.41 25.35 98.59 73.24 53.23 29.84 45.97 71.77 Latent Optimization 50.00 53.12 54.69 40.62 52.11 35.21 47.89 63.38 39.52 37.90 60.48 61.29
Figure 3: Visual analysis on text-guided protein editing with latent optimization. (a-c) visualize structure editing with more/less $\alpha$-helices editing. (d-f) visualize structure editing with more/less $\beta$-sheets editing. (g-i) visualize peptide binding editing on PDB 3IQI.
Downstream Task: Protein Property Prediction
Last but not least, we assess the intermediate protein representation learned from the contrastive pretraining step using the ProteinCLAP model. ProteinCLAP aligns the text-protein sequence pairs, and the protein modality uses an encoder initialized from ProtBERT [6]. ProtBERT was pretrained on BFD in a masked language model (MLM) manner [27, 28], a self-supervised learning method to predict the masked tokens, and only the protein sequence information is included. We compare the original pretrained ProtBERT with the updated representation from ProteinCLAP.
Experiment setting. We consider the following six tasks gathered from TAPE [48]. Secondary structure (SS) prediction [49] is a token-level prediction task, and the goal is to predict the secondary structure (helix, sheet, etc.) for each amino acid. SS-Q3 and SS-Q8 refer to the three-way and eight-way classifications, respectively. Homology prediction [50, 51] is to predict the protein fold of the protein sequence. The evaluation metric is accuracy for these three tasks. Contact [52, 53, 44] is to predict if pairs of amino acids within a protein sequence pair are in contact ($<8\accentset{\circ}{A}$ apart), and the metric is the precision for the top $L/2$ contacts, where $L$ is the length of the medium- and long-range proteins. Fluorescence [54] is to predict the log-fluorescence intensity of a protein. Stability [40] is to predict a proxy for protein stability. Both tasks are measured using Spearman’s $\rho$ metric.
Baselines. LSTM [55] and ResNet [56] are the two classic representation methods for sequential data. TAPE Transformer [48] is a Transformer specific for protein sequence representation. MSA Transformer [5] adds the multiple sequence alignments (MSAs) into the Transformer. The most comparable models to ProteinDT are ProtBERT [6] and its variant because ProteinCLAP shares the same backbone model with ProtBERT. OntoProtein [57] is a variant of ProtBERT and should be highlighted because (1) it also starts the pretraining from a pretrained ProtBERT, and (2) it utilizes the same modalities for pretraining, i.e., textual description and protein sequence. Yet the difference is that OntoProtein conducts a contrastive pretraining based on a knowledge graph (KG). The KG covers 565K proteins, 47K Gene Ontology (GO) terms [21], and 31 relations.
Observations. The property prediction results are in Table 3. We show two variants of ProteinCLAP pretraining, corresponding to the EBM-NCE loss and InfoNCE loss, respectively. We can observe that both ProteinCLAP pretrained representations can consistently reach the best performance on three structure-related prediction tasks. Yet for the other two tasks, Homology and Stability, they are still within the top three models. The only exception is the Fluorescence task, where all the ProtBERT models are performing worse than other baselines.
Table 3: Results on six protein property prediction tasks from TAPE. The best results and the second-best results are marked. Baseline results are replicated from OntoProtein, with the exception of ProtBERT and OntoProtein for Contact prediction, both of which were rerun.
Method Structure Evolutionary Engineering SS-Q3 $\uparrow$ SS-Q8 $\uparrow$ Contact $\uparrow$ Homology $\uparrow$ Fluorescence $\uparrow$ Stability $\uparrow$ LSTM 0.75 0.59 0.26 0.26 0.67 0.69 TAPE Transformer 0.73 0.59 0.25 0.21 0.68 0.73 ResNet 0.75 0.58 0.25 0.17 0.21 0.73 MSA Transformer - 0.73 0.49 - - ProtBERT 0.81 0.67 0.59 0.29 0.61 0.82 OntoProtein 0.82 0.68 0.56 0.24 0.66 0.75 ProteinDT-ProteinCLAP-InfoNCE (Ours) 0.8331 0.6935 0.6031 0.3092 0.6029 0.8130 ProteinDT-ProteinCLAP-EBM-NCE (Ours) 0.8334 0.6919 0.6007 0.2841 0.6105 0.7952
Discussion
Challenge regarding dataset quality. In this work, we construct a text-protein pair dataset called ProteinDT. It has 441K data pairs, yet such a dataset size is small compared to the vision and language domains. If we want to take the experimental protein backbone structure into consideration, that would be reduced to merely 45K pairs of data [58]. Thus, data insufficiency has become the bottleneck of this research direction. ProteinDT solves this by adopting pretrained checkpoints, while more promising solutions exist. For example, it may be possible to create alternative datasets from full-text biomedical manuscripts by using named entity recognition to identify proteins and mapping their names to identifiers and sequences [59]. This approach may generate a larger, more diverse, but lower-quality text corpus paired with a more restricted set of proteins.
Challenge regarding evaluation. For the text-guided protein design tasks, the textual description is quite flexible, but the output proteins need to be tested. Although most protein properties can be tested experimentally, the cost and throughput of the experiments can vary considerably depending on the property. In this work, we take the computationally feasible solutions with regressor models as surrogate oracles, following the evaluation approach of reinforcement learning models [60]. However, the reliability of such surrogate models is not guaranteed. To mitigate this dependency, we incorporate alternative evaluation methods, such as ESMFold and GNINA, during the editing assessment. Although the preliminary protein editing results are quite encouraging, additional evaluations will be required to more thoroughly assess the strengths and weaknesses of ProteinDT on diverse editing tasks.
Outlook. One ambitious vision of ProteinDT is to enable zero-shot protein engineering for specific protein functions. Deep learning models can guide protein design when trained on large high-throughput experimental datasets [61, 62], which are only available for a few protein functions. Specialized approaches reduce the data requirements but still require some function-specific experimental data [63, 8]. General protein language models can have a good correlation with specific functional assays in a zero-shot setting but cannot distinguish different properties of the same protein [64, 7]. ProteinDT shows a potential strategy for jump-starting future protein design of specific functions without requiring experimental data by leveraging the rich information in the textual description.
Methods
Backgrounds on Generative Models
Autoregressive Model
The autoregressive (AR) model is one of the state-of-the-art generative paradigms for estimating the distribution of discrete sequential data [65, 66, 67], including text sequences and protein sequences. Suppose the data can be formulated as a sequence of $N$ components as ${\bm{x}}=({\bm{x}}^{(1)},{\bm{x}}^{(2)},\dots,{\bm{x}}^{(N)})$, such as amino acids in a protein sequence. The corresponding data distribution is factorized in an autoregressive manner: $p({\bm{x}})=p({\bm{x}}^{(1)}|{\bm{x}}^{(0)})p({\bm{x}}^{(2)}|{\bm{x}}^{(0:1)})\dots p({\bm{x}}^{(N)}|{\bm{x}}^{(0:N-1)})$, and we set ${\bm{x}}^{(0)}=\emptyset$ and $p({\bm{x}}^{(1)}|{\bm{x}}^{(0)})=p({\bm{x}}^{(1)})$ as the starting points. Parameterized by $p_{\theta}({\bm{x}}^{(n)}|{\bm{x}}^{(0:n-1)}),n\in\{1,2,\dots N\}$, the training objective is next-token prediction:
$\mathcal{L}=\mathbb{E}_{{\bm{x}}}\big{[}-\sum_{n=1}^{N}\log p({\bm{x}}^{(n)}|{\bm{x}}^{(0:n-1)})\big{]}.$ (1)
During the inference stage, running the AR model repeatedly can generate the sequence token-by-token.
Denoising Diffusion Generative Model
The key idea for the diffusion generative model is to gradually map the data to random noise, and generate samples by the parameterized denoising process. In specific, given a data point (e.g., protein sequence) from the data distribution ${\bm{x}}_{0}\sim q({\bm{x}}_{0})$, the forward process generates a Markov chain of latent variables ${\bm{x}}_{1},{\bm{x}}_{2},\ldots,{\bm{x}}_{T}$ via the following conditional distributions:
$q({\bm{x}}_{t}|{\bm{x}}_{t-1})=\mathcal{N}({\bm{x}}_{t};\sqrt{1-\beta_{t}}{\bm{x}}_{t-1},(\beta_{t})I),$ (2)
where $\{\beta_{t}\}_{t=1}^{T}$ is the variance schedule, and it keeps non-decreasing as $t\to T$. Such a forward process can be viewed as progressively adding noise to ground-truth data ${\bm{x}}_{0}$, and finally nearly leads to a complete noise ${\bm{x}}_{T}\sim p(x_{T})$.
The reverse process aims to gradually denoise a randomly sampled noise ${\bm{x}}_{T}$, so as to generate the given data point ${\bm{x}}_{0}$. To attain this goal,  [68] derives a simplified surrogate objective function, as follows:
$\mathcal{L}=\mathbb{E}_{t\sim[1,T],{\bm{x}}_{0}\sim p({\bm{x}}_{0})}\big{[}\|\epsilon_{\theta}({\bm{x}}_{t},t)-\epsilon\|^{2}\big{]},$ (3)
where $\epsilon\sim\mathcal{N}(0,I)$ is drawn from a Gaussian noise distribution and $\epsilon_{\theta}({\bm{x}}_{t},t)$ is a parameterized neural network learning to estimate the noise. The above equation is equivalent to the denoising score matching method [69, 70], and both can be united under a stochastic differential equation (SDE) framework [71]. In this work, we are interested in the discrete version of the diffusion model and we implement ProteinDT under the SDE framework to benefit future exploration.
For notation clarity, in what follows, we use ${\bm{x}}_{\text{t}}$ to represent the protein sequence data and $\tilde{\bm{x}}_{t}$ to represent the diffused data at timestep $t$ in the diffusion model.
ProteinCLAP Model
Maximizing the mutual information (MI) between modalities has been widely used for representation pretraining [72, 73]. The intuition is that MI measures the non-linear dependency between random variables (e.g., text and protein), and thus maximizing MI between modalities can force them to share more critical information. Along this research line, contrastive learning has proven its effectiveness in various tasks [74, 75, 11, 76], and the high-level idea is to transform the density estimation problem into a classification problem. Thus, we adopt it to align the representation of two modalities, coined as Contrastive LAnguage and Protein pretraining (ProteinCLAP).
To train ProteinCLAP, first, we extract positive and negative pairs from SwissProtCLAP: a text-protein pair is labeled as positive if and only if it is extracted from the same protein record in SwissProtCLAP. Then, ProteinCLAP aligns the positive pairs by increasing their similarity scores and contrasts the negative pairs by decreasing their similarity scores. In ProteinCLAP, the similarity score (e.g., cosine similarity) can be easily calculated by mapping the text and protein sequence to the representation space. The last stage is to conduct the alignment and contrasting simultaneously, where we consider two commonly-used loss functions, EBM-NCE [76] and InfoNCE [74]:
$\mathcal{L}_{\text{EBM-NCE}}=-\frac{1}{2}\mathbb{E}_{{\bm{x}}_{\text{p}},{\bm{x}}_{\text{t}},{\bm{x}}_{\text{t}}^{\prime}}\big{[}\log\sigma(E({\bm{x}}_{\text{p}},{\bm{x}}_{\text{t}})+\log(1-\sigma(E({\bm{x}}_{\text{p}},{\bm{x}}_{\text{t}}^{\prime}))\big{]}-\frac{1}{2}\mathbb{E}_{{\bm{x}}_{\text{p}},{\bm{x}}_{\text{p}}^{\prime},{\bm{x}}_{\text{t}}}\big{[}\log\sigma(E({\bm{x}}_{\text{p}},{\bm{x}}_{\text{t}})+\log(1-\sigma(E({\bm{x}}_{\text{p}}^{\prime},{\bm{x}}_{\text{t}}))\big{]},$ (4)
$\mathcal{L}_{\text{InfoNCE}}=-\frac{1}{2}\mathbb{E}_{{\bm{x}}_{\text{p}},{\bm{x}}_{\text{t}}}\Big{[}\log\frac{\exp(E({\bm{x}}_{\text{p}},{\bm{x}}_{\text{t}}))}{\exp(E({\bm{x}}_{\text{p}},{\bm{x}}_{\text{t}}))+\sum_{{\bm{x}}_{\text{t}}^{\prime}}\exp(E({\bm{x}}_{\text{p}},{\bm{x}}_{\text{t}}^{\prime}))}\Big{]}-\frac{1}{2}\mathbb{E}_{{\bm{x}}_{\text{p}},{\bm{x}}_{\text{t}}}\Big{[}\log\frac{\exp(E({\bm{x}}_{\text{p}},{\bm{x}}_{\text{t}}))}{\exp(E({\bm{x}}_{\text{p}},{\bm{x}}_{\text{t}}))+\sum_{{\bm{x}}_{\text{p}}^{\prime}}\exp(E({\bm{x}}_{\text{p}}^{\prime},{\bm{x}}_{\text{t}}))}\Big{]},$ (5)
where ${\bm{x}}_{\text{t}}^{\prime}$ and ${\bm{x}}_{\text{p}}^{\prime}$ are negative samples randomly sampled from the empirical data distribution, $E(\cdot,\cdot)$ is an energy (scalar) function defined over two modalities and we take the dot product of the representation accordingly [77], i.e., $E({\bm{x}}_{\text{t}},{\bm{x}}_{\text{p}})=\langle{\bm{z}}_{\text{t}},{\bm{z}}_{\text{p}}\rangle$. The difference between EBM-NCE (Equation 4) and InfoNCE (Equation 5) is that the former performs binary classification and the latter conducts multiclass classification. Although it has been shown that EBM-NCE is superior on the graph-related contrastive pretraining tasks [78, 76, 79], here we consider both of them for a comprehensive hyperparameter search progress.
By doing so, ProteinCLAP forms informative representations for proteins, and such robust representations help improve representation-oriented downstream tasks such as protein property prediction, as discussed in the Results Section. Another benefit of ProteinCLAP is its mapping of representations between textual descriptions and protein sequences. As a result, such a multi-modal bridge empowers retrieving protein representations from textual descriptions, and such mapping is further leveraged by a facilitator model in ProteinDT to facilitate text-guided protein generation and editing. We will discuss the facilitator component in detail next.
ProteinFacilitator Model
The ProteinFacilitator model in ProteinDT aims at mapping a piece of text (a text prompt describing the properties of the target protein) to a protein representation that captures both the semantic information of the text and essential protein sequence patterns. This protein representation is then subsequently used for protein generation and editing. Such an explicit representation mapping from text modality to protein modality is inspired by the foundation models developed in vision [13] and chemistry [18], which has shown to improve the generalization quality (e.g., image diversity).
Concretely, the ProteinFacilitator model explicitly produces the protein sequence representation from the text representation, and we use a Gaussian distribution to model such a process. Note that this is just a simplifying choice of assumption and not a limitation of our method. We have text-to-protein generation and text-guided protein editing results in the Results Section (Table 1-2, Figure 3) verifying that such an assumption can give us reasonably good results.
We introduce an alignment module $f_{\text{ProteinFacilitator}}(\cdot)$ and then we estimate the ProteinFacilitator distribution using a conditional Gaussian as $p({\bm{z}}_{\text{p}}|{\bm{z}}_{\text{t}})=\mathcal{N}({\bm{z}}_{\text{p}};f_{\text{ProteinFacilitator}}({\bm{z}}_{\text{t}}),I)$. To minimize its likelihood, the objective is essentially the representation reconstruction (regression) task [76, 18]:
$\mathcal{L}_{\text{ProteinFacilitator}}=\|f_{\text{ProteinFacilitator}}({\bm{z}}_{\text{t}})-{\bm{z}}_{\text{p}})\|^{2},$ (6)
where $f_{\text{ProteinFacilitator}}(\cdot)$ is parameterized by a multi-layer perception. The produced protein sequence representation, i.e., $\tilde{\bm{z}}_{\text{p}}=f_{\text{ProteinFacilitator}}({\bm{z}}_{\text{t}})$, will be treated as the input condition, i.e., ${\bm{z}}_{\text{c}}=\tilde{\bm{z}}_{\text{p}}$, for the decoder model, as will be discussed in the next section. It is worth noting that, we also directly use the text sequence representation encoded by ProteinCLAP as the condition, namely ${\bm{z}}_{\text{c}}={\bm{z}}_{\text{t}}$. In the Results Section, we empirically verify the effectiveness of introducing the ProteinFacilitator model on the text-to-protein design task.
Protein Decoder Model
With the condition ${\bm{z}}_{\text{c}}$ generated by the aforementioned ProteinFacilitator, a decoder model is applied to conduct conditional generative modeling for protein generation, i.e., $p({\bm{x}}_{\text{p}}|{\bm{z}}_{\text{c}})$. In detail, during the training stage of ProteinDT, we directly take ${\bm{z}}_{\text{c}}={\bm{z}}_{\text{p}}$ to train the decoder in a teacher-forcing manner. In the next two sections, we introduce two types of conditional generative models for protein sequence generation: autoregressive (AR) and diffusion generative (ProteinDiff).
An Autoregressive Decoder
The training objective of AR is the next-token prediction (in Equation 1). As for model architecture, we adopt the Transformer for AR modeling [26], which is one of the most expressive architectures for various AR tasks, such as text-to-text [66, 67], text-to-speech [80, 81], and text-to-image generation [82]. More concretely, here we train a T5-Base model from scratch [67]. For conditional protein generation, the T5 decoder will generate the protein sequence token-by-token, where each generated token is conditioned on both the input condition and the previously decoded tokens. An illustration is shown in Supplementary C.
A Discrete Diffusion Decoder
We additionally investigate the protein sequence decoder employing the diffusion model within the ProteinDT framework. The diffusion generative model has been prevalent to solve the density estimation of vision tasks[70, 68, 83, 71]. More recently, the multinomial diffusion model [84, 85] has been proposed for modeling and generating discrete data such as textual sequence. We adopt the multinomial diffusion for protein sequence generation and call it ProteinDiff. The key component of ProteinDiff consists of a forward process that systematically introduces noise to the protein sequence ${\bm{x}}_{\text{p}}$ and a reverse process that performs denoising to reconstruct the original protein sequence. An illustration is shown in Supplementary C.
Forward process. This process adds noise to the input at each of the $T$ timesteps to convert $\tilde{\bm{x}}_{0}\to\tilde{\bm{x}}_{T}$, where $\tilde{\bm{x}}_{0}={\bm{x}}_{\text{p}}$ is the protein sequence and $\tilde{\bm{x}}_{T}$ is a random noise following a predefined distribution as will be introduced next. Specifically, given a condition representation ${\bm{z}}_{\text{c}}$, the forward transition of ProteinDiff at timestep $t$ is defined as:
$q(\tilde{\bm{x}}_{t}|\tilde{\bm{x}}_{t-1},{\bm{z}}_{\text{c}})=\text{Categorical}(\tilde{\bm{x}}_{t};p=\tilde{\bm{x}}_{t-1}Q_{t}),$ (7)
where $\tilde{\bm{x}}_{t}$ is the diffused protein sequence and $Q_{t}$ is the transition matrix at timestep $t$, respectively. We further take the absorbing transition [85] when adding noises: $Q_{t}=(1-\beta_{t})I+\beta\mathbbm{1}e_{m}^{T}$, where $\beta_{t}$ is the variance scheduled at timestep $t$ and $e_{m}$ is a one-hot vector with one on the absorbing state $m$ and zeros elsewhere. Specifically, we take 30 tokens as the vocabulary (including 26 alphabetic characters and 4 special tokens such as “[PAD]”), and the “[MASK]” token is used as the absorbing state $m$. Other variations of the transition matrix $Q_{t}$ are reserved for future exploration [85].
Objective and parameterization. To parameterize the above diffusion process, we adopt the following training tricks. Recent works [85, 86, 87] have empirically shown that the $\tilde{\bm{x}}_{0}$-paramtererization benefits more for the discrete data in diffusion modeling with reduced stochasticity. We therefore adopt such a strategy. That is, at each timestep $t$, we train a conditional transition network $p_{\theta}(\tilde{\bm{x}}_{0}|\tilde{\bm{x}}_{t},{\bm{z}}_{\text{c}})$ to predict the ground-truth protein sequence tokens. By taking the expectation over all $T$ timesteps, the objective function of ProteinDiff is the following for the masked tokens:
$\mathcal{L}_{\text{ProteinDiff}}=\mathbb{E}_{\tilde{\bm{x}}_{0}}\mathbb{E}_{t}\mathbb{E}_{\tilde{\bm{x}}_{t}|\tilde{\bm{x}}_{0}}\big{[}-\log p_{\theta}\big{(}\tilde{\bm{x}}_{0}|\tilde{\bm{x}}_{t},{\bm{z}}_{\text{c}}\big{)}\big{]}.$ (8)
The aforementioned transition network takes as input the condition representation ${\bm{z}}_{\text{c}}$ and the noised sequence $\tilde{\bm{x}}_{t}$ at timestep $t$, and outputs the predicted masked amino acids at timestep $t+1$. We consider two transition network architectures, a recurrent neural network (RNN) transition network and a Transformer transition network (BERT-Base) [25], for modeling.
Reverse process and decoding. The reverse process $\tilde{\bm{x}}_{T}\to\tilde{\bm{x}}_{0}$ in the diffusion model corresponds to the protein sequence sampling for protein generation in ProteinDiff. In detail, to generate a protein sequence, we first generate a diffused sequence $\tilde{\bm{x}}_{T}$ with all tokens being the absorbing state $m$. Then we consider two sampling strategies. (1) The simplified sampling [87]: for timestep $t$, we sample the mask positions and predict the masked tokens following the distribution $p_{\theta}(\tilde{\bm{x}}_{0}|\tilde{\bm{x}}_{t+1},{\bm{z}}_{\text{c}})$ for $\tilde{\bm{x}}_{t-1}$. (2) The weighted sampling [85]: for timestep $t$, the $p_{\theta}(\tilde{\bm{x}}_{t}|\tilde{\bm{x}}_{t+1})$ is indeed $p_{\theta}(\tilde{\bm{x}}_{0}|\tilde{\bm{x}}_{t+1},{\bm{z}}_{\text{c}})$ weighted by a posterior distribution (more details are in Supplementary C). We repeat such sampling process for $T$ timesteps, from $\tilde{\bm{x}}_{T}$ to $\tilde{\bm{x}}_{0}$, which yields protein sequence generation.
Code Availability
The source code can be found at this GitHub repository or Zenodo [88].
Data Availability
The dataset generation scripts can be found at this GitHub repository. We also provide the preprocessed pretraining dataset (SwissProtCLAP) at this HuggingFace link.
Acknowledgements
This project was partly done during Shengchao Liu’s internship at Nvidia and Ph.D. program at Mila-UdeM, and was supported in part by the Natural Sciences and Engineering Research Council (NSERC) Discovery Grant, the Canada CIFAR AI Chair Program, collaboration grants between Microsoft Research and Mila, Samsung Electronics Co., Ltd., Amazon Faculty Research Award, Tencent AI Lab Rhino-Bird Gift Fund, two NRC Collaborative R&D Projects, IVADO Fundamental Research Project grant PRF-2019-3583139727, and NSF award CHE 2226451.